OpenStudy (anonymous):

can someone help me with the factor of 6m^2+4m-16 i don't want the answer I want someone to help me threw the steps and help me if I mess up please

OpenStudy (anonymous):

Use quadratic formula x = -b + or - the square toot of b^2 - 4ac all over 2a werer a=6 b=4 and c=16

OpenStudy (asnaseer):

There is a little known method called the slide and divide method that you might find interesting. See here: http://algebrafunsheets.com/blog/2008/11/22/slide-and-divide-method-of-factoring-trinomials/ Let me know if you need any help in understanding that method.

OpenStudy (anonymous):

or you can factor find your factors of 6 and your factors of 16 the sum or difference of your factors should equal 4

OpenStudy (anonymous):

ok im stuck on this

OpenStudy (asnaseer):

have you looked at the link I gave above?

OpenStudy (anonymous):

yes ok so ac is 6*1 and b is 4

OpenStudy (asnaseer):

where did you get that from?

OpenStudy (asnaseer):

the equation you are trying to factor is:\[6m^2+4m-16\]correct?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

so, if you read that link, the first thing it asks you to do is: "(Always ensure to pull out any common factors before continuing with this method)" Can you see a common factor across all terms in your equation?

OpenStudy (anonymous):

the m and maybe 4 and 16

OpenStudy (asnaseer):

"m" does not divide into ALL the terms - so it cannot be a common factor same goes for 4 and 16 find a factor that divides into ALL the terms

OpenStudy (asnaseer):

what divides into all these terms: \(6m^2, 4m\) and \(16\)

OpenStudy (asnaseer):

hint: 6, 4 and 16 are all EVEN numbers so they they can ALL be divided by ?

OpenStudy (anonymous):

2

OpenStudy (asnaseer):

great!

OpenStudy (asnaseer):

so now we can first rewrite your equation as:\[6m^2+4m−16=2(3m^2+2m-8)\]agreed?

OpenStudy (anonymous):

yes i see how you did that

OpenStudy (asnaseer):

ok good, so now we can just concentrate of factoring this:\[3m^2+2m−8\]

OpenStudy (asnaseer):

now the slide and divide method says that we first take the coefficient of \(m^2\) and "slide" it over to the "constant" term using multiplication

OpenStudy (asnaseer):

so here we need to remove the 3 from \(3m^2\) and "slide" it over to the "-8" replacing "-8" with "-8"*3 = "-24"

OpenStudy (asnaseer):

that will leave us with:\[m^2+2m-24\]agreed?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

ok, the next step is to then try to factor this expression - which should be a lot simpler. we need to find the ?'s in this expression: \(m^2+2m-24=\)(m ?)(m ?)

OpenStudy (asnaseer):

so, can you think of two number that when multiplied together give you 24?

OpenStudy (asnaseer):

i.e. find all the factors of 24

OpenStudy (asnaseer):

you would then pick out those two factors that can also be combined (i.e. added or subtracted) to give you 2 (for the \(2m\) term)

OpenStudy (asnaseer):

e.g. a) 24 = 1 x 24 but 24 and 1 cannot be combined to give you 2 - so we reject these two b) 24 = 2 x 12 but " ... " etc

OpenStudy (anonymous):

24= 3 x 8 24=4 x 6

OpenStudy (asnaseer):

good - so which of those give you factors that can be combined to give you 2?

OpenStudy (asnaseer):

combined as in added or subtracted

OpenStudy (anonymous):

4 and 6

OpenStudy (asnaseer):

perfect - and how can we get 2 by combining 4 and 6?

OpenStudy (anonymous):

-4 + 6

OpenStudy (asnaseer):

great! so now we can write:\[m^2+2m-24=(m-4)(m+6)\]agreed?

OpenStudy (anonymous):

yes thats what i got :)

OpenStudy (asnaseer):

perfect! :) now the next step is to divide the constants in this expression by the number we "slid" over at the beginning - in this case we slid over 3, so we need to divide the constants by 3 and then simplify to get:\[(m-\frac{4}{3})(m+\frac{6}{3})=(m-\frac{4}{3})(m+2)\]do you follow?

OpenStudy (anonymous):

not really

OpenStudy (asnaseer):

do you remember we "slid" over the 3 at the beginning? recall we started with:\[3m^2+4m-16\]

OpenStudy (anonymous):

yes o ok so now we use the 3

OpenStudy (asnaseer):

sorry I mean\[3m^2+2m-8\]

OpenStudy (asnaseer):

let me list the steps from the start just to clarify for you...

OpenStudy (anonymous):

ok

OpenStudy (asnaseer):

\[\begin{align} 6m^2+4m-16&=2(3m^2+2m-8)\\ \text{next we slide over the 3 to get:}\\ m^2+2m-24\\ \text{we then factorised this to get:}\\ (m-4)(m+6)\\ \text{we now divide by 3 - the number we slid over:}\\ (m-\frac{4}{3})(m-2) \end{align}\]

OpenStudy (asnaseer):

the lines shown above from line 2 onwards are intermediate steps we need to do in order to factorise \(3m^2+2m-8\)

OpenStudy (anonymous):

ok

OpenStudy (asnaseer):

the final step says that if you get left with any fractions (as we have), then you take the denominator of the fraction and move it over to the "m" term, so the term:\[m-\frac{4}{3}\]becomes:\[(3m-4)\]we can then conclude that:\[3m^2+2m-8=(3m-4)(m+2)\]NOTE: I made a mistake on my last line above - (m-2) term should have been (m+2)

OpenStudy (asnaseer):

putting it all together we get:\[6m^2+4m-16=2(3m^2+2m-8)=2(3m-4)(m+2)\]hope that makes sense.

OpenStudy (asnaseer):

I would advise you to re-read that link I gave above and try some more examples of your own to become familiar with this method.

OpenStudy (asnaseer):

the best way to generate examples is to start from what you want the answer to be and expand it - then see if your factorisation process leads to the same answer. e.g. \[(2x+3)(x-5)=2x^2-7x-15\]so see if you can use this method to factorise \(2x^2-7x-15\)

OpenStudy (asnaseer):

ping me if you need more help in understanding this - and good luck! :)

OpenStudy (anonymous):

thanks so much and i will

OpenStudy (asnaseer):

ok - and keep at it my friend - practice makes perfect! :)

OpenStudy (anonymous):

ok im stuck on the one that you wanted me to try i think i am doing it right but im not shore

OpenStudy (asnaseer):

please list your steps so I can then try to spot where you may have made a mistake

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2x^2-7x+30 2x^2-7x+30 = (x+10) X-3) 10m-2 = 5

OpenStudy (asnaseer):

it should be "2x^2 - 7x - 15" at the start

OpenStudy (anonymous):

o yea i did have that down i typed it wrong sorry

OpenStudy (anonymous):

ok let me re type it

OpenStudy (asnaseer):

you also had a plus sign instead of a minus sign

OpenStudy (anonymous):

2x^2 - 7x - 15 x^2-7x+30 (x+10) (x-3) 10x-2= 5

OpenStudy (asnaseer):

you've made the same mistake here -> the second line should have a "- 30" not a "+ 30"

OpenStudy (asnaseer):

let me do this one for you and I'll give you another one to try...

OpenStudy (asnaseer):